A grocery store may increase the original price of a product to cover the expenses of running the store.The markup is the amount that is added to the original cost to create the selling price (the price the consumer pays)of the groceries . The percent markup is called the markup rate and is usually expressed as a percent of the original price .Taking the original cost and adding the markup calculates the selling price of an item. The formula S = C + rC can be used to help find the selling price. The selling price,S is equal to the original cost, C,plus the markup rate,r,multiplied by the original cost, C.
A 20-lb turkey has a 35% markup and is selling for a price of $22.50. Use the selling price formula to find the original cost of the turkey. Round to the nearest cent.
This is what I need to solve.
Thank you
Dorian

Answer by scott8148(6628)

solve the formula for C and plug in the given values

factoring ___ S = C(1 + r)

dividing ___ S / (1 + r) = C

22.50 / (1 + .35) = C

22.50 / 1.35 = C